### v1

# Quantum conditional mutual information as a probe of measurement-induced entanglement phase transitions

Abstract
We propose that the quantum conditional mutual information (QCMI), computed with a suitably chosen partition of the system, serves as a powerful probe for detecting measurement-induced entanglement phase transitions in monitored quantum circuits. To demonstrate this, we investigate monitored variable-range Clifford circuits and identify the phase boundary between volume-law and area-law entanglement phases by performing finite-size scaling analyses of the QCMI. Assuming that the entanglement entropy exhibits a logarithmic dependence on system size at criticality in short-range interacting cases, we further show that the QCMI allows for the simultaneous determination of both the critical point and the universal coefficient of the logarithmic term in the entanglement entropy via a crossing-point analysis. For the shortest-range interacting case studied, we obtain the thermodynamic-limit value of the coefficient as $\tilde{c}=1.519(3)$, which is significantly smaller than values reported in previous studies.
I Introduction
Measurement-induced phase transitions (MIPTs) in quantum circuits composed of unitary dynamics and local projective measurements have attracted significant interest in recent years [1, 2, 3, 4, 5]. These transitions separate two distinct entanglement phases: a volume-law phase, in which the entanglement entropy of a subsystem scales proportionally with its volume, and an area-law phase, where it scales with the boundary area of the subsystem. Such transitions emerge from the competition between unitary dynamics, which generally generate entanglement among qubits, and local projective measurements, which tend to disentangle the measured qubits from the rest of the system. Remarkably, measurement-induced entanglement phase transitions [6, 7], as well as other nonequilibrium phase transitions [8], have been demonstrated experimentally on real quantum devices. These experimental breakthroughs have been made possible either by employing mid-circuit measurement and reset functionalities [7, 8], or by using a reference qubit initially maximally entangled with the monitored system as a probe for the entanglement phase transition [6, 9].
To numerically detect the measurement-induced entanglement phase transitions, the entanglement entropy itself is not always a suitable quantity, particularly due to its emergent logarithmic dependence on subsystem size at criticality in (1+1)$d$ systems [4, 2]. Indeed, performing finite-size scaling with entanglement entropy requires subtracting its value at an estimated critical point (e.g., a critical measurement rate) to eliminate the logarithmic dependence [4], which introduces an additional source of uncertainty. As an alternative, the tripartite mutual information (TMI) [10], defined as $I_{3}(A,B,C)=S_{A}+S_{B}+S_{C}-S_{A\cup B}-S_{B\cup C}-S_{A\cup C}-S_{A\cup B\cup C}$, where $S_{X}$ is the entanglement entropy of subsystem $X$, has been used to prove MIPTs [9, 11]. With an appropriately chosen tripartiton $A,B,$ and $C$, this linear combination of entanglement entropies effectively distinguishes volume-law and area-law phases without being affected by the logarithmic dependence at criticality. This approach has proven successful in systems with variable-range (power-law decaying) interactions [12, 13, 14, 15] as well as in systems with sparse nonlocal interactions [16]. While the effectiveness of the TMI is empirically established, it remains unclear–at least to our knowledge–why the logarithmic contribution is suppressed, especially given that $I_{3}(A,B,C)$ involves an odd number (seven) of entanglement entropy terms.
In this paper, we propose that another linear combination of entanglement entropies, of the form [17],
$I(A:C|B)=S_{A\cup B}+S_{B\cup C}-S_{B}-S_{A\cup B\cup C},$ (1)
with an appropriately chosen partition $A,B,$ and $C$ of the system, serves as an effective and pertinent probe of MIPTs. The quantity $I(A:C|B)$ is known as the quantum conditional mutual information (QCMI), and has been recently employed to quantify entanglement growth in stabilizer circuits composed of nearest-neighbor two-qubit Clifford gates [18]. Here, we demonstrate that the QCMI is also well-suited for probing measurement-induced entanglement phase transitions, through numerical simulations of one-dimensional variable-range ($1/r^{\alpha}$) Clifford circuits interspersed with local projective measurements. While similar models have been studied previously [13, 15], our circuit definition differs slightly in detail. Through careful finite-size-scaling analyses of $I(A:C|B)$, we find that no phase transition occurs for $\alpha\lesssim 2$, whereas for $2\lesssim\alpha\lesssim 3$, measurement-induced entanglement phase transitions emerge with critical exponents distinct from those in nearest-neighbor circuits and in the regime $3\lesssim\alpha$. These findings are qualitatively consistent with earlier studies on monitored circuits with variable-range interactions, not restricted to Clifford circuits [12, 13, 14, 15]. Furthermore, using crossing-point analysis of $I(A,C|B)$, we extract the coefficient $\tilde{c}$ of the emergent logarithmic scaling of the entanglement entropy at criticality in the thermodynamic limit. For the shortest-range interacting case studied, we find $\tilde{c}=1.519(3)$, which is significantly smaller than previous estimates obtained in monitored nearest-neighbor random Clifford circuits [2, 11, 19].
The rest of this paper is organized as follows. In Sec. II, we define the stabilizer circuit with variable-range two-qubit Clifford gates and illustrate how the QCMI behaves at criticality under a suitable partition of the system. In Sec. III, we numerically demonstrate that the QCMI not only detects the measurement-induced entanglement transition, but also enables estimation of the coefficient of the logarithmic term in the entanglement entropy at criticality. In Sec. IV, we discuss the correlation-length critical exponent and the universal coefficient of the logarithmic term, and compare them with values obtained in previous studies. Finally, we present conclusions and remarks in Sec. V.
II Model and method
II.1 Stabilizer circuit
We investigate entanglement phase transitions in stabilizer circuits composed of variable-range two-qubit Clifford gates and local projective measurements. Stabilizer circuits are well suited for our purposes, as they enable efficient simulation of the dynamics of thousands of qubits, not only in weakly entangled states but also in highly entangled ones. This stands in contrast to exact diagonalization or tensor network state approaches, which are typically limited to systems with smaller system sizes or low entanglement.
We consider a one-dimensional chain of $L$ qubits with periodic boundary conditions, assuming that $L$ is even. Let $|\phi_{\zeta}(t)\rangle$ denote the pure quantum state of the entire system of $L$ qubits at discrete time step (simply referred to as “time”) $t$, initialized as $|\phi_{\zeta}(t=0)\rangle=|0\rangle^{\otimes L}$. Here, $\zeta$ indexes different circuit realizations characterized by random seeds. Time $t$ is incremented by one after each application of $L$ two-qubit Clifford gates, followed by projective measurements on each qubit along the Pauli-$Z$ axis with probability $p$ (see dashed horizontal lines in Fig. 1). Each two-qubit Clifford gate is randomly selected from the complete set of 11520 distinct two-qubit Clifford gates [20, 21], and applied to a randomly chosen pair of qubits $i$ and $j(\not{=}i)$ with probability
$P(r_{ij})=\frac{1}{\cal N}\frac{1}{r_{ij}^{\alpha}},$ (2)
where $r_{ij}$ is the distance between qubits $i$ and $j$ under periodic boundary conditions, defined as $r_{ij}=|i-j|$ for $|i-j|\leqslant L/2$ and $r_{ij}=L-|i-j|$ for $|i-j|>L/2$, ${\cal N}=\sum_{x=1}^{L/2}\frac{1}{x^{\alpha}}$ is the normalization constant, and $\alpha$ is a parameter that controls the interaction range of the two-qubit gates. We simulate circuit dynamics using the tableau algorithm [22] (see also Supplemental Material [23]), for system sizes ranging from $L=16$ to $2048$, and evolution times up to $t_{\rm max}=4096$. The parameters $\alpha$ and $p$ are varied in the ranges $1\leqslant\alpha\leqslant 4$ and $10\leqslant p<$ 1, respectively.
II.2 Entanglement entropy and quantum conditional mutual information
At each time step $t$, we compute the von Neumann entanglement entropies of the state $|\phi_{\zeta}(t)\rangle$ for subsystems $X=A\cup B,B\cup C,B$, and $A\cup B\cup C$ (see Fig. 2) using the base-2 logarithm:
$S_{X,\zeta}(t)=-{\rm Tr}_{X}\left[\hat{\rho}_{X,\zeta}(t)\log_{2}\hat{\rho}_{X,\zeta}(t)\right],$ (3)
where $\hat{\rho}_{X}(t)={\rm Tr}_{\overline{X}}\left[|\phi_{\zeta}(t)\rangle\langle\phi_{\zeta}(t)|\right]$ is the reduced density matrix of subsystem $X$, and $\bar{X}$ denotes its complement. To compute the entanglement entropy in stabilizer circuits, we follow the method described in Ref. [24] (see also Supplemental Material [23]). Note that when $p=1$, the entanglement entropy vanishes for all subsystems since the state $|\phi_{\zeta}(t)\rangle$ becomes a product state immediately after every measurement.
Figure 1: Schematic figure of a monitored variable-range random Clifford circuit. Time flows from bottom to top. Vertical lines represent qubits, dumbbell-shaped symbols indicate two-qubit random Clifford gates, and circles denote projective measurements that occur probabilistically.
After sufficiently long time evolution, the state $|\phi_{\zeta}(t)\rangle$ reaches a steady state in which the entanglement entropy $S_{X,\zeta}(t)$ fluctuates around a saturation value. Such a steady state can exhibit a phase transition between volume-law and area-law phases as the parameter $p$ or $\alpha$ is varied, in the thermodynamic limit ($L\to\infty$). From the entanglement entropy $S_{X,\zeta}(t)$, we evaluate the steady-state QCMI as
$I(A:C|B)=\frac{1}{N_{t}}\sum_{k=0}^{N_{t}-1}\left\langle\left\langle I_{\zeta}(t_{\rm min}+k\Delta_{I,t})\right\rangle\right\rangle,$ (4)
where $I_{\zeta}(t)\equiv S_{A\cup B,\zeta}(t)+S_{B\cup C,\zeta}(t)-S_{B,\zeta}(t)-S_{A\cup B\cup C,\zeta}(t)$ is the QCMI for the state $|\phi_{\zeta}(t)\rangle$. Here, $t_{\rm min}$ is the time after which $I_{\zeta}(t)$ has reached saturation on average, $\Delta_{I,t}$ is the sampling time interval, $N_{t}$ is the number of time steps used in the time average, and $\langle\langle\cdots\rangle\rangle\equiv\frac{1}{R}\sum_{\zeta=1}^{R}\cdots$ denotes the statistical average over $R$ independent circuit realizations. We set $t_{\rm min}=\min(2L,2048)$ and $R=10560$, and determine $\Delta_{I,t}$ as the estimated autocorrelation time of $\left\langle\left\langle I_{\zeta}(t)\right\rangle\right\rangle$ for $t\geqslant t_{\rm min}$, assuming exponential decay of the corresponding autocorrelation function. In the worst case, near criticality for $L=2048$, $\Delta_{I,t}\sim 100$, while it is significantly smaller away from criticality. $N_{t}$ is chosen as the maximum integer satisfying $t_{\rm min}+(N_{t}-1)\Delta_{I,t}\leqslant t_{\rm max}$.
II.3 Entanglement entropy at criticality
Motivated by the emergent conformal criticality observed in monitored Clifford and non-Clifford circuits with short-range interactions or hoppings [4, 2, 25, 26, 12, 13], we assume that the steady-state entanglement entropy $S_{X}=\frac{1}{N_{t}}\sum_{k=0}^{N_{t}-1}\langle\langle S_{X,\zeta}(t_{\rm min}+k\Delta_{S_{X},t})\rangle\rangle$ at criticality behaves as [27]
$S_{X}\sim\frac{c}{3}\log_{2}{l_{X}}+c^{\prime}\quad(p=p_{\rm c}),$ (5)
where $l_{X}\equiv\frac{L}{\pi}\sin\frac{\pi|X|}{L}$ is the chord length of subsystem $X$ with $|X|$ denoting the number of qubits in $X$, $p_{\rm c}$ the critical value of the measurement rate, $c$ a universal coefficient of the logarithmic term (often referred to as an effective central charge) [28], and $c^{\prime}$ a nonuniversal constant. The autocorrelation time $\Delta_{S_{X},t}$ for the steady-state entanglement entropy is estimated in the same manner as $\Delta_{I,t}$, and exhibits similar behavior. Notice that the coefficient $c$ here is not equivalent to the central charge in conformal field theory. For example, in the case of the percolation universality class, the central charge is known to be zero [28, 29]. Note also that in the deep volume-law phase ($p\ll p_{c}$), the entanglement entropy scales extensively as $S_{X}\sim|X|$, while in the area-law phase ($p\gg p_{\rm c}$), it scales with the boundary, $S_{X}\sim|\partial X|$, where $|\partial X|$ denotes the number of qubits on the boundary of subsystem $X$.
Figure 2: Two distinct partitionings of the $L$-qubit system into subsystems $A$ (green), $B$ (white), $C$ (red), and $\overline{A\cup B\cup C}$ (gray). Each circle represents a qubit, and the number of qubits in each segment is indicated. Partition (a) is used for the QCMI in Eq. (7), while partition (b) is employed for the QCMI in Eq. (8).
II.4 Quantum conditional mutual information at criticality
To use the QCMI as a probe for the measurement-induced entanglement phase transition, we partition the system according to the following three conditions (see also Fig. 2): (i) The system is divided into four parts $A,B,C,$ and $\overline{A\cup B\cup C}$, where $A,B,$ and $C$ are contiguous segments located in the middle of the periodic chain. This setup ensures that each entanglement entropy $S_{X}$ is associated with a different chord length $l_{X}$, while all share the same nonuniversal constant $c^{\prime}$. By substituting Eq. (5) into the definition of QCMI in Eq. (1), we obtain
$I(A:C|B)\sim\frac{c}{3}\log_{2}\frac{l_{A\cup B}l_{B\cup C}}{l_{B}l_{A\cup B\cup C}}\quad(p=p_{\rm c}),$ (6)
indicating that the nonuniversal constant $c^{\prime}$ cancels out in $I(A:C|B)$ at criticality. (ii) To eliminate the dependence of $I(A:C|B)$ on the total system size $L$ at criticality, we choose the subsystem sizes such that $|A\cup B|$, $|B\cup C|$, $|B|$, and $|A\cup B\cup C|$ are all proportional to $L$. (iii) To capture the volume-law entanglement phase, the subsystem sizes are chosen such that $L_{A\cup B}+L_{B\cup C}\neq L_{B}+L_{A\cup B\cup C}$, because for $p\ll p_{\rm c}$, the QCMI behaves as $I(A:C|B)\sim L_{A\cup B}+L_{B\cup C}-L_{B}-L_{A\cup B\cup C}$ , where $L_{X}\equiv\min(|X|,L-|X|)$ is the smaller of the number of qubits in $X$ and its complement $\overline{X}$. We note that in the area-law phase $p\gg p_{\rm c}$, the QCMI behaves as $I(A:C|B)\sim 0$, regardless of the partitioning. Moreover, the inequality $I(A:C|B)\geqslant 0$ always holds due to the strong subadditivity of von Neumann entanglement entropy [30]. Thus, as long as the partitioning satisfies conditions (i)-(iii), the QCMI $I(A:C|B)$ serves as a meaningful probe of the measurement-induced entanglement phase transition. Importantly, condition (iii) is independent of the assumption of conformal criticality in Eq. (5), implying that $I(A:C|B)$ can detect the entanglement phase transition even in the absence of emergent conformal invariance.
Specifically, we consider two types of partitioning for the same system, in a spirit of Ref. [31]. The first partitioning sets $|A|=|B|=|C|=L/4$ [see Fig. 2(a)], for which we find
$I_{a}(A:C|B)\sim\frac{c}{3}\log_{2}{\frac{\sin^{2}{\frac{\pi}{2}}}{\sin{\frac{\pi}{4}}\sin{\frac{3\pi}{4}}}}=\frac{c}{3}\quad\quad(p=p_{\rm c}).$ (7)
The second partitioning is given by $|A|=|C|=\frac{5L}{16}$ and $|B|=\frac{L}{4}$ [see Fig. 2(b)], for which we obtain
$I_{b}(A:C|B)\sim\frac{c}{3}\log_{2}{\frac{\sin^{2}{\frac{9\pi}{16}}}{\sin{\frac{\pi}{4}}\sin{\frac{\pi}{8}}}}\approx 1.83\times\frac{c}{3}\quad\quad(p=p_{\rm c}).$ (8)
Here, the subscripts $a$ and $b$ in $I_{a}$ and $I_{b}$ correspond to the partitionings illustrated in Figs. 2(a) and 2(b), respectively. It should be emphasized that Eqs. (6)–(8) are valid only under the assumption of logarithmic scaling of the entanglement entropy at criticality, as expressed in Eq. (5). As we show in the next section, our numerical results strongly support the validity of this assumption for $\alpha\gtrsim 3$.
III Results
III.1 Phase diagram
Before presenting a detailed analysis, we first show the global phase diagram of the stabilizer circuits with variable-range two-qubit Clifford gates as a function of $\alpha$ and $p$ in Fig. 3. Using the QCMI as a probe, we accurately determine the phase boundary separating the area-law and volume-law entanglement phases for $\alpha>2$. In what follows, we analyze the critical properties of the system along this phase boundary. Additionally, we confirm that no area-law phase exists for $\alpha\lesssim 2$, which is consistent with the previous analytical arguments in the high-measurement-rate regime [13, 15].
Figure 3: Phase diagram in the thermodynamic limit. The phase boundary is estimated from a data-collapse analysis of the QCMI $I_{a}(p,L)$, evaluated using the partitioning scheme shown in Fig. 2(a). An equivalent phase diagram is obtained using the QCMI $I_{b}(p,L)$ based on the alternative partitioning shown in Fig. 2(b). The dashed vertical line at $\alpha=2$ is a guide to the eye.
III.2 Quantum conditional mutual information near criticality
Let us denote the QCMI as $I_{\gamma}(p,L)$, where $\gamma=a$ and $b$ correspond to the two distinct partitioning schemes shown in Figs. 2(a) and 2(b), respectively. At the critical point $p=p_{\mathrm{c}}$, the QCMI becomes independent of the system size [see Eqs. (7) and (8)], which suggests that the crossing-point analysis [32] can be effectively used to estimate both $p_{\mathrm{c}}$ and the correlation-length critical exponent $\nu$. Indeed, as shown in Fig. 4, the curves of $I_{\gamma}(p,L)$ for different $L$ appear to cross at a single point for sufficiently large system sizes. The estimated critical points $p_{\mathrm{c}}$ are nearly identical for both partitionings $\gamma=a$ and $b$, for each value of $\alpha$. This already demonstrates, at least in part, that the QCMI serves as a reliable probe to detect the critical point. For smaller values of $\alpha$, deviations from the size-independent behavior are more pronounced in smaller systems [see Figs. 4(b) and 4(d)]. Nevertheless, well-established methods for handling corrections to scaling have been developed and documented in the literature [33]. Interestingly, we find that corrections to scaling becomes more subtle for larger $\alpha$. As discussed in Sec. III.4, nonmonotonic finite-size behavior of the crossing points [see Figs. 8(a) and 9(a)] emerges in the vicinity of the critical point, suggesting that multiple correction terms may need to be considered. A similar issue was only recently resolved in the study of the Néel-paramagnetic quantum phase transition in dimerized Heisenberg models, where the known value of the correlation-length exponent $\nu$ from the expected O(3) universality class was crucial for the analysis [34]. In our case, however, neither the critical exponent $\nu$ nor the universality class is known a priori, necessitating the use of an alternative approach, as described in the following.
Figure 4: QCMI $I_{\gamma}(p,L)$ for different partitionings and interaction ranges. Panels (a) and (b) show results for $\gamma=a$, while (c) and (d) correspond to $\gamma=b$. The interaction-range parameter is $\alpha=3.5$ for (a) and (c), and $\alpha=2.5$ for (b) and (d). Vertical dashed lines indicate the critical points estimated from the data-collapse analyses in Figs 6 and 7.
III.3 Data-collapse
Here, we adopt an empirical approach to estimate $p_{\mathrm{c}}$ and $\nu$ in the thermodynamic limit based on a data-collapse analysis. We employ the Bayesian scaling analysis method [35] to collapse the data of $I_{\gamma}(p,L)$, assuming the following scaling ansatz:
$I_{\gamma}(p,L)=f_{\gamma}(\Delta pL^{1/\nu}),$ (9)
where $\Delta p=p-p_{\mathrm{c}}$ is the deviation from the critical point, and $f_{\gamma}$ denotes the scaling function. Representative examples of successful data-collapse fits are shown in Fig. 5. We do not explicitly include a correction term of the form $\lambda L^{-\omega}$ in the scaling function, where $\lambda$ is an irrelevant scaling field and $\omega$ is the associated correction exponent. Instead, we perform data-collapse fits over the system size range $L\in[L_{\mathrm{min}},L_{\mathrm{max}}]$ with $L_{\mathrm{max}}=2048$ fixed and $L_{\mathrm{min}}$ systematically varied. We then examine the asymptotic behavior of the estimated values of $p_{\mathrm{c}}$ and $\nu$ as $L_{\mathrm{min}}$ increases [36, 37]. An important aspect of this analysis is the estimation of uncertainties in the fitted parameters. Because data collapsing is inherently a nonlinear procedure, the results are sensitive to the choice of initial parameters in the Bayesian analysis as well as the selected data set. To mitigate these sensitivities and obtain reliable error bars, we apply the bootstrap resampling method. The resulting uncertainties of the estimated $p_{\mathrm{c}}$ and $\nu$ are shown in Figs. 6 and 7. We extrapolate the results to the thermodynamic limit, i.e., $01/L_{\mathrm{min}}=$ 0, using a linear fit based on the four largest values of $L_{\mathrm{min}}$. Although the use of a linear fit and the specific number of data points are empirical choices, we find that the estimated values of $p_{\mathrm{c}}$ and $\nu$ in the thermodynamic limit are consistent between $\gamma=a$ and $\gamma=b$ within two standard deviations. As summarized in Table 1, the results obtained for other values of $\alpha$ are overall consistent between the two different partitioning schemes, agreeing to at least four (three) significant digits for $p_{\mathrm{c}}$ ($\nu$) in most cases of $\alpha$. Additional analyses are provided in Supplemental Material [23].
Figure 5: Data-collapse fits of the QCMI $I_{\gamma}(p,L)$. Panels (a) and (b) show results for $\gamma=a$, while (c) and (d) correspond to $\gamma=b$. The interaction-range parameter is $\alpha=3.5$ for (a) and (c), and $\alpha=2.5$ for (b) and (d). The extrapolated values of $p_{\mathrm{c}}$ and $\nu$ in the thermodynamic limit, as obtained in Figs 6 and 7, are used here.
Figure 6: Extrapolation of (a) $p_{\mathrm{c}}$ and (b) $\nu$ obtained from data-collapse analysis using the QCMI $I_{\gamma}(p,L)$ for $L\geqslant L_{\mathrm{min}}$ at $\alpha=3.5$. Solid and dashed lines represent linear fits to the data for $\gamma=a$ and $\gamma=b$, respectively. The extrapolated values in the thermodynamic limit ($1/L_{\mathrm{min}}\to 0$) are indicated by solid symbols at $01/L_{\mathrm{min}}=$ 0.
Figure 7: Same as Fig. 6, but for $\alpha=2.5$.
Table 1: Summary of the critical point $p_{\mathrm{c}}$ and the correlation-lemgth exponent $\nu$ in the thermodynamic limit, estimated from data-collapse analysis of the QCMI $I_{a}$ and $I_{b}$ for the two different partitioning schemes. The numbers in parentheses indicate the uncertainty in the last digits.
$p_{\mathrm{c}}$ $\nu$
$\alpha$ $I=I_{a}$ $I=I_{b}$ $I=I_{a}$ $I=I_{b}$
4.0 0.20363(1) 0.203617(9) 1.255(4) 1.248(4)
3.8 0.21007(1) 0.210075(9) 1.253(6) 1.242(5)
3.6 0.218241(9) 0.218239(9) 1.257(5) 1.249(5)
3.5 0.22320(1) 0.22318(1) 1.249(6) 1.239(5)
3.4 0.228815(7) 0.228803(7) 1.245(5) 1.240(4)
3.2 0.24262(1) 0.24261(1) 1.249(6) 1.247(5)
3.1 0.25113(2) 0.25110(2) 1.246(8) 1.249(8)
3.0 0.26089(1) 0.26086(1) 1.259(5) 1.256(7)
2.9 0.27219(2) 0.27215(2) 1.270(8) 1.270(6)
2.85 0.27847(1) 0.27844(1) 1.286(6) 1.280(6)
2.8 0.28522(1) 0.28519(1) 1.304(6) 1.299(7)
2.75 0.29248(2) 0.29244(1) 1.323(7) 1.316(6)
2.7 0.30025(1) 0.30020(1) 1.333(6) 1.332(5)
2.6 0.317526(8) 0.317473(8) 1.384(4) 1.388(4)
2.5 0.33726(1) 0.33723(1) 1.455(6) 1.457(6)
2.4 0.35963(2) 0.35964(2) 1.54(1) 1.54(1)
2.3 0.38508(7) 0.38526(8) 1.74(2) 1.78(2)
2.2 0.4135(2) 0.4140(2) 2.0(1) 2.1(1)
2.1 0.4500(7) 0.4521(8) 3.2(2) 3.4(2)
2.0 0.502(2) 0.514(6) 13(5) 13(4)
III.4 Crossing-point analysis
With the accurate estimates of $p_{\mathrm{c}}$ and $\nu$ obtained from the data-collapse analysis, we now return to the crossing-point analysis. Crossing points are defined as the values of $p$ and $I_{\gamma}$ at which the curves of $I_{\gamma}(p,L)$ and $I_{\gamma}(p,2L)$ intersect. These points are denoted as $p_{\gamma}^{\times}(L,2L)$ and $I_{\gamma}^{\times}(L,2L)$, respectively. A key difficult in this analysis is the nonmonotonic finite-size behavior of the crossing points, as clearly illustrated in Figs. 8(a) and 9(a). To capture this behavior, it is necessary to consider at least two correction terms in the scaling forms:
$p_{\gamma}^{\times}(L,2L)=p_{\mathrm{c}}+L^{-1/\nu}\left(a_{1}L^{-\omega_{1}}+a_{2}L^{-\omega_{2}}\right)$ (10)
and
$I_{\gamma}^{\times}(L,2L)=\frac{c}{3}+b_{1}L^{-\omega_{1}}+b_{2}L^{-\omega_{2}},$ (11)
where $a_{i}$ and $b_{i}$ ($i$=1,2) are constants, and the correction exponents are ordered such that $\omega_{2}>\omega_{1}>0$. Note that $I_{b}^{\times}(L,2L)$ is normalized by a constant factor ($\simeq 1.83$) so that its thermodynamic limit matches that of $I_{a}^{\times}(L,2L)$, i.e., $c/3$, [see Eqs. (7) and (8)]. There are two possible scenarios for the origin of the second correction term [34]. One possibility is that it arises from a second irrelevant scaling field, independent of the first. The other is that it corresponds to the second higher-order term of the same irrelevant filed as the first one, i.e, $\omega_{2}=2\omega_{1}$. Although the scaling forms in Eqs. (10) and (11) appear relatively simple, we find that fitting both correction terms with sufficient precision is quite challenging [34, 33]. This difficulty persists even when we fix $p_{\mathrm{c}}$ and $\nu$ in Eq. (10) to the values obtained from the data-collapse analysis in Table 1.
As a result, we are unable to reliably test the first scenario involving two independent correction exponents by treating $\omega_{1}$ and $\omega_{2}$ as free parameters. Instead, motivated by the second scenario, we achieve reasonable fits by fixing $p_{\mathrm{c}}$ and $\nu$, and imposing $\omega_{2}=2\omega_{1}$ in Eq. (10), as shown in Fig. 8. It is worth noting that the fitted values of $\omega_{1}$ are consistent between $\gamma=a$ and $\gamma=b$ (see Table 2 for various values of $\alpha$). This consistency strongly supports the second scenario. Moreover, the obtained values of $\omega_{1}$ are compatible with the cross-point data of $I_{\gamma}^{\times}(L,2L)$ when fitted to Eq. (11), where only $c$ and $b_{i}$ are fitting parameters, as shown in Fig. 9. We also find that the obtained values of $c$ are consistent between the two different partitioning schemes at least for $3.0\lesssim\alpha$ (see Table 2), as will be discussed in the next section. This further supports the assumption $\omega_{2}=2\omega_{1}$.
Figure 8: System-size dependence of the crossing points $p_{\gamma}^{\times}(L,2L)$ for (a) $\alpha=3.5$ and (b) $\alpha=2.5$. Open circles (triangles) represent the results for $\gamma=a$ ($b$), while solid ones at $01/L=$ 0 indicate the values of $p_{\mathrm{c}}$ obtained by the data-collapse analysis shown in Figs. 6 and 7, as well as in Table 1. Solid and dashed lines are independent fits to the data for the two partitioning schemes, $\gamma=a$ and $\gamma=b$, respectively, using the functional from given in Eq. (10), where $p_{\mathrm{c}}$ and $\nu$ are fixed at the extrapolated values in the thermodynamic limit obtained by the data-collapse analysis, and $\omega_{2}=2\omega_{1}$ is assumed. The fitting values of $\omega_{1}$ are indicated in each panel.
Figure 9: System-size dependence of the QCMI at the crossing points for (a) $\alpha=3.5$ and (b) $\alpha=2.5$. Open circles (triangles) represent the results for $\gamma=a$ ($b$). Solid and dashed lines are fits to the data using the functional form given in Eq. (11), where $\omega_{1}$ is fixed to the values obtained from the fits of $p_{\gamma}^{\times}(L,2L)$ in Fig. 8, and $\omega_{2}=2\omega_{1}$ is assumed.
Table 2: Summary of the correction exponent $\omega_{1}$ and the coefficient $c/3$ obtained from the crossing-point analysis of $p_{\gamma}^{\times}(L,2L)$ and $I_{\gamma}^{\times}(L,2L)$, assuming $\omega_{2}=2\omega_{1}$. The numbers in parentheses indicate the uncertainty in the last digits.
$\omega_{1}$ $c/3$
$\alpha$ $I=I_{a}$ $I=I_{b}$ $I=I_{a}$ $I=I_{b}$
4.0 0.94(5) 0.92(7) 1.053(2) 1.052(2)
3.8 0.93(4) 0.85(5) 1.050(3) 1.044(3)
3.6 0.94(3) 0.93(4) 1.058(1) 1.058(2)
3.5 0.88(2) 0.88(3) 1.055(2) 1.058(2)
3.4 0.90(1) 0.85(3) 1.061(1) 1.061(2)
3.2 0.83(2) 0.80(4) 1.056(1) 1.059(2)
3.1 0.76(2) 0.76(4) 1.049(2) 1.057(2)
3.0 0.76(2) 0.75(3) 1.048(1) 1.059(3)
2.9 0.70(2) 0.70(3) 1.034(2) 1.053(3)
2.85 0.73(2) 0.68(2) 1.035(2) 1.050(3)
2.8 0.68(2) 0.67(3) 1.021(3) 1.044(4)
2.75 0.67(2) 0.65(3) 1.010(4) 1.039(5)
2.7 0.65(2) 0.67(3) 1.004(4) 1.044(5)
2.6 0.61(2) 0.60(3) 0.984(4) 1.030(5)
2.5 0.57(2) 0.57(2) 0.971(5) 1.034(5)
2.4 0.55(1) 0.53(1) 0.987(7) 1.060(8)
2.3 0.51(1) 0.46(1) 1.025(8) 1.123(9)
2.2 0.52(1) 0.44(2) 1.093(7) 1.24(1)
2.1 0.43(2) 0.41(2) 1.228(8) 1.394(8)
2.0 0.55(3) 0.8(2) 1.01(1) 0.84(2)
IV Discussion
In Fig. 10, we plot the correlation-length exponent $\nu$ estimated from the data-collapse analysis (see Table 1) as a function of $\alpha$. We confirm that for $\alpha\gtrsim 3$, our model belongs to the same universality class as that of the short-range interacting models [13, 15]. In particular, our numerical estimates are in full agreement with recent results for the model with nearest-neighbor Clifford gates [19], and differ significantly from the value predicted by classical two-dimensional percolation theory, $\nu=4/3$.
The existence of a universality class behavior for $\alpha\gtrsim 3$ is further supported by the coefficient $c$ of the logarithmic term in the entanglement entropy, as shown in Fig. 11. The close agreement of $c$ values estimated from the two different partitioning schemes, $\gamma=a$ and $b$, strongly supports the presence of emergent conformal criticality as described by Eqs. (5)–(8) for $\alpha\gtrsim 3$, in line with the behavior observed in the short-range interacting models. In contrast, the noticeable discrepancy in the estimated values of $c$ between the two partitioning schemes for $\alpha\lesssim 3$ indicates a breakdown of the scaling relations in Eqs. (5)–(8). This deviation from emergent conformal criticality corroborates previous findings of non-conformal-field-theory criticality, as evidenced by the deviation of the dynamical critical exponent $z$ from unity in Ref. [13]. It is worth noting, however, that the critical point $p_{c}$ and the exponent $\nu$ remain consistent between the two partitioning schemes for all values of $\alpha$ (see Table 1). This consistency highlights that the QCMI remains a valid and robust probe of the measurement-induced entanglement phase transition, even in the absence of emergent conformal criticality.
Figure 10: $\alpha$-dependence of the correlation-length critical exponent $\nu$. The dotted line with the gray shaded area represents the result of the nearest-neighbor model, $\nu=1.265(15)$ [19]. The inset shows a magnified view near $\alpha=3.0$. For comparison, the prediction of classical two-dimensional percolation theory, $\nu=4/3$, is also indicated by the dashed line.
Now we compare the coefficient $c/3$ of the logarithmic term in the entanglement entropy with those obtained for monitored nearest-neighbor random Clifford circuits in Refs. [2, 11, 19]. While our analysis uses the base-2 logarithm, previous studies employed the natural logarithm in defining this coefficient. To enable direct comparison, we convert our result as
$\tilde{c}=\frac{1}{\ln 2}\times\frac{c}{3}\simeq 1.443\times\frac{c}{3},$ (12)
so that $\tilde{c}$ is consistent with the convention used in the earlier works. From our estimate of $c/3=1.053(2)$ for the shortest-range case $\alpha=4$ using the partitioning scheme $\gamma=a$ (see Table 2), we obtain $\tilde{c}=1.519(3)$. This value is noticeably smaller than those in the previous studies: $\tilde{c}=1.6$ with up to 512 qubits [2], $\tilde{c}=1.61(3)$ with up to 512 qubits [11], and $\tilde{c}=1.57(1)$ with up to 10240 qubits [19] for monitored nearest-neighbor random Clifford circuits. Furthermore, we find that $c/3$ exhibits only minor variation across different partitioning schemes ($\gamma=a,b$) and for a range of $\alpha\gtrsim 3$ [see the inset of Fig. 11 and Table 2]. This supports the conclusion that our estimate of $\tilde{c}$ remains consistently smaller than those in the previous works, even when accounting for uncertainties.
Figure 11: Coefficient $c/3$ of the logarithmic term in the entanglement entropy as a function of $\alpha$. Circles and triangles represent the results for $\gamma=a$ and $\gamma=b$, respectively. The inset shows a magnified view of the region $3\leqslant\alpha\leqslant 4$.
To clarify the origin of this discrepancy, we also estimate $\tilde{c}$ from the entanglement entropy at criticality, as shown in Fig. 12. Using the critical point $p_{\mathrm{c}}$ obtained from the data-collapse analysis, we fit the entanglement entropy of the form given in Eq. (5), treating $c/3$ and $c^{\prime}$ as fitting parameters. We obtain $\tilde{c}=1.6126(7)$, $1.5484(9)$, and $1.5263(7)$ [$c/3=1.1175(5)$, $1.0730(6)$, and $1.0577(5)$] for $\alpha=3.0$, $3.5$, and $4.0$, respectively. These results indicate that $c/3$ obtained from the entanglement entropy approaches the value estimated from the QCMI from above as $\alpha$ increases (see Supplemental Material [23] for results at other values of $\alpha$). Since finite-size effects are expected to be smaller at larger $\alpha$, the estimates of $\tilde{c}$ from the entanglement entropy at criticality are more reliable for larger $\alpha$ than for the smaller ones. Nevertheless, we argue that the estimates of $\tilde{c}$ based on the QCMI are generally more accurate, as the QCMI enables crossing-point analyses by eliminating the logarithmic dependence at criticality, as shown in Eqs. (7) and (8). Indeed, the values of $\tilde{c}$ estimated via the QCMI exhibit much weaker dependence on $\alpha$ [see the inset of Fig. 11 and Table 2] compared to those obtained from the entanglement entropy [23]. We therefore adopt $\tilde{c}=1.519(3)$, obtained from the QCMI for $\alpha=4.0$, as our final estimate for the shortest-range interacting case.
Figure 12: Entanglement entropy at criticality as a function of chord length for (a) $\alpha=4.0$, (b) $\alpha=3.5$, and (c) $\alpha=3.0$. Results are shown for various subsystems, $X=A\cup B$, $B\cup C$, and $D=\overline{A\cup B\cup C}$, under partitioning schemes $\gamma=a$ and $b$, using different symbols and colors. Solid lines represent fits to the data using the form given in Eq. (5). The estimated values of $c/3$ are also indicated in each panel.
V Conclusion
Using stabilizer circuits with variable-range two-qubit Clifford gates and local projective measurements, we have demonstrated that the QCMI, $I(A:C|B)$, serves as a powerful probe of the measurement-induced entanglement phase transition. An important advantage of the QCMI is its ability to not only distinguish between volume-law and area-law entanglement phases, but also to provide access to the universal coefficient $c$ of the logarithmic term in the entanglement entropy at criticality [see Eq. (6)]. Through careful finite-size scaling analyses, we have determined the phase diagram and obtained reliable estimates of both the correlation-length critical exponent $\nu$ and the universal coefficient $c$. Notably, for short-range interacting cases ($\alpha\gtrsim 3$), the consistency of the estimated values of $c$ across two distinct partitioning schemes, each yielding different values of $\frac{1}{3}\log_{2}\frac{l_{A\cup B}l_{B\cup C}}{l_{B}l_{A\cup B\cup C}}$ in Eq. (6), further validates the robustness of our approach. Finally, the finite-size scaling analysis scheme based on the QCMI presented in this work is not limited to stabilizer circuits. Its extension to more general, non-stabilizer circuits remains an important direction for future research.
Acknowledgements.
A portion of the numerical simulations was performed using the HOKUSAI supercomputer at RIKEN (Project ID: Q22525) and the supercomputer Fugaku installed at RIKEN R-CCS (Project ID: hp220138). This work was supported by Grant-in-Aid for Scientific Research (C) (No. JP18K03475, No. JP21K03395, No. JP22K03520, and No. JP24K06894) and Grant-in-Aid for Scientific Research (A) (No. JP21H04446) from MEXT, Japan. Additional support was provided by the UTokyo Quantum Initiative, the RIKEN TRIP initiative (RIKEN Quantum), and the COE research grant in computational science from Hyogo Prefecture and Kobe City through Foundation for Computational Science. We also acknowledge funding from JST COI-NEXT (Grant No. JPMJPF2221) and the Program for Promoting Research on the Supercomputer Fugaku (Grant No. MXP1020230411) from MEXT, Japan.

### v2

@@
 I Introduction
-Measurement-induced phase transitions (MIPTs) in quantum circuits composed of unitary dynamics and local projective measurements have attracted significant interest in recent years [1, 2, 3, 4, 5].
+Measurement-induced phase transitions (MIPTs) in quantum circuits composed of unitary dynamics and local projective measurements have attracted significant interest in recent years [Li-PRB2018, Li-PRB2019a, Chan-PRB2019, Skinner-PRX2019, Fisher-ARCMP2023].
 These transitions separate two distinct entanglement phases: a volume-law phase, in which the entanglement entropy of a subsystem scales proportionally with its volume, and an area-law phase, where it scales with the boundary area of the subsystem.
 Such transitions emerge from the competition between unitary dynamics, which generally generate entanglement among qubits, and local projective measurements, which tend to disentangle the measured qubits from the rest of the system.
-Remarkably, measurement-induced entanglement phase transitions [6, 7], as well as other nonequilibrium phase transitions [8], have been demonstrated experimentally on real quantum devices.
-These experimental breakthroughs have been made possible either by employing mid-circuit measurement and reset functionalities [7, 8], or by using a reference qubit initially maximally entangled with the monitored system as a probe for the entanglement phase transition [6, 9].
-To numerically detect the measurement-induced entanglement phase transitions, the entanglement entropy itself is not always a suitable quantity, particularly due to its emergent logarithmic dependence on subsystem size at criticality in (1+1)$d$ systems [4, 2].
-Indeed, performing finite-size scaling with entanglement entropy requires subtracting its value at an estimated critical point (e.g., a critical measurement rate) to eliminate the logarithmic dependence [4], which introduces an additional source of uncertainty.
-As an alternative, the tripartite mutual information (TMI) [10], defined as $I_{3}(A,B,C)=S_{A}+S_{B}+S_{C}-S_{A\cup B}-S_{B\cup C}-S_{A\cup C}-S_{A\cup B\cup C}$, where $S_{X}$ is the entanglement entropy of subsystem $X$, has been used to prove MIPTs [9, 11].
+Remarkably, measurement-induced entanglement phase transitions [Noel-NatPhys2022, Koh-NP2023], as well as other nonequilibrium phase transitions [Chertkov_NatPhys2023], have been demonstrated experimentally on real quantum devices.
+These experimental breakthroughs have been made possible either by employing mid-circuit measurement and reset functionalities [Koh-NP2023, Chertkov_NatPhys2023], or by using a reference qubit initially maximally entangled with the monitored system as a probe for the entanglement phase transition [Noel-NatPhys2022, Gullans-PRX2020].
+To numerically detect the measurement-induced entanglement phase transitions, the entanglement entropy itself is not always a suitable quantity, particularly due to its emergent logarithmic dependence on subsystem size at criticality in (1+1)$d$ systems [Skinner-PRX2019, Li-PRB2019a].
+Indeed, performing finite-size scaling with entanglement entropy requires subtracting its value at an estimated critical point (e.g., a critical measurement rate) to eliminate the logarithmic dependence [Skinner-PRX2019], which introduces an additional source of uncertainty.
+As an alternative, the tripartite mutual information (TMI) [Kitaev-PRL2006], defined as $I_{3}(A,B,C)=S_{A}+S_{B}+S_{C}-S_{A\cup B}-S_{B\cup C}-S_{A\cup C}-S_{A\cup B\cup C}$, where $S_{X}$ is the entanglement entropy of subsystem $X$, has been used to prove MIPTs [Gullans-PRX2020, Zabalo-PRB2020].
 With an appropriately chosen tripartiton $A,B,$ and $C$, this linear combination of entanglement entropies effectively distinguishes volume-law and area-law phases without being affected by the logarithmic dependence at criticality.
-This approach has proven successful in systems with variable-range (power-law decaying) interactions [12, 13, 14, 15] as well as in systems with sparse nonlocal interactions [16].
+This approach has proven successful in systems with variable-range (power-law decaying) interactions [Minato-PRL2022, Block-PRL2022, Muller-PRL2022, Sahu-PRB2022] as well as in systems with sparse nonlocal interactions [Hashizume-PRR2022].
 While the effectiveness of the TMI is empirically established, it remains unclear–at least to our knowledge–why the logarithmic contribution is suppressed, especially given that $I_{3}(A,B,C)$ involves an odd number (seven) of entanglement entropy terms.
-In this paper, we propose that another linear combination of entanglement entropies, of the form [17],
+In this paper, we propose that another linear combination of entanglement entropies, of the form [Levin-PRL2006],
 $I(A:C|B)=S_{A\cup B}+S_{B\cup C}-S_{B}-S_{A\cup B\cup C},$ (1)
 with an appropriately chosen partition $A,B,$ and $C$ of the system, serves as an effective and pertinent probe of MIPTs.
-The quantity $I(A:C|B)$ is known as the quantum conditional mutual information (QCMI), and has been recently employed to quantify entanglement growth in stabilizer circuits composed of nearest-neighbor two-qubit Clifford gates [18].
+The quantity $I(A:C|B)$ is known as the quantum conditional mutual information (QCMI) [Cerf-PRL1997, Cerf-PhysicaD1998], and has been recently employed to quantify measurement-induced topological phase transitions [Lavasani-NatPhys2021, Lavasani-PRL2021, Kells-SciRep2023], and entanglement growth in stabilizer circuits composed of nearest-neighbor two-qubit Clifford gates [Sang-PRXQ2023].
+The QCMI has also been used to characterize one-dimensional topological insulators and superconductors at equilibrium [Maiellaro-PRR2022, Maiellaro-PRB2023].
 Here, we demonstrate that the QCMI is also well-suited for probing measurement-induced entanglement phase transitions, through numerical simulations of one-dimensional variable-range ($1/r^{\alpha}$) Clifford circuits interspersed with local projective measurements.
-While similar models have been studied previously [13, 15], our circuit definition differs slightly in detail.
+While similar models have been studied previously [Block-PRL2022, Sahu-PRB2022], our circuit definition differs slightly in detail.
+We show that the QCMI is free from multiplicative and additive renormalization; at criticality, it becomes system-size independent under suitable partitioning of the system.
+This property makes the QCMI particularly useful for identifying critical points and extracting universal quantities much like the Binder ratio in conventional statistical physics.
+Notably, the value of the QCMI at emergent conformal criticality directly yields the universal coefficient of the logarithmic term.
 Through careful finite-size-scaling analyses of $I(A:C|B)$, we find that no phase transition occurs for $\alpha\lesssim 2$, whereas for $2\lesssim\alpha\lesssim 3$, measurement-induced entanglement phase transitions emerge with critical exponents distinct from those in nearest-neighbor circuits and in the regime $3\lesssim\alpha$.
-These findings are qualitatively consistent with earlier studies on monitored circuits with variable-range interactions, not restricted to Clifford circuits [12, 13, 14, 15].
+These findings are qualitatively consistent with earlier studies on monitored circuits with variable-range interactions, not restricted to Clifford circuits [Minato-PRL2022, Block-PRL2022, Muller-PRL2022, Sahu-PRB2022].
 Furthermore, using crossing-point analysis of $I(A,C|B)$, we extract the coefficient $\tilde{c}$ of the emergent logarithmic scaling of the entanglement entropy at criticality in the thermodynamic limit.
-For the shortest-range interacting case studied, we find $\tilde{c}=1.519(3)$, which is significantly smaller than previous estimates obtained in monitored nearest-neighbor random Clifford circuits [2, 11, 19].
+For the shortest-range interacting case studied, we find $\tilde{c}=1.519(3)$, which is significantly smaller than previous estimates obtained in monitored nearest-neighbor random Clifford circuits [Li-PRB2019a, Zabalo-PRB2020, Sierant-PRB2022].
 The rest of this paper is organized as follows.
@@
 1).
-Each two-qubit Clifford gate is randomly selected from the complete set of 11520 distinct two-qubit Clifford gates [20, 21], and applied to a randomly chosen pair of qubits $i$ and $j(\not{=}i)$ with probability
+Each two-qubit Clifford gate is randomly selected from the complete set of 11520 distinct two-qubit Clifford gates [Corcoles-PRA2013, Richter-PRR2023], and applied to a randomly chosen pair of qubits $i$ and $j(\not{=}i)$ with probability
 $P(r_{ij})=\frac{1}{\cal N}\frac{1}{r_{ij}^{\alpha}},$ (2)
-where $r_{ij}$ is the distance between qubits $i$ and $j$ under periodic boundary conditions, defined as $r_{ij}=|i-j|$ for $|i-j|\leqslant L/2$ and $r_{ij}=L-|i-j|$ for $|i-j|>L/2$, ${\cal N}=\sum_{x=1}^{L/2}\frac{1}{x^{\alpha}}$ is the normalization constant, and $\alpha$ is a parameter that controls the interaction range of the two-qubit gates.
-We simulate circuit dynamics using the tableau algorithm [22] (see also Supplemental Material [23]), for system sizes ranging from $L=16$ to $2048$, and evolution times up to $t_{\rm max}=4096$.
-The parameters $\alpha$ and $p$ are varied in the ranges $1\leqslant\alpha\leqslant 4$ and $10\leqslant p<$ 1, respectively.
+where $r_{ij}$ is the distance between qubits $i$ and $j$ under periodic boundary conditions, defined as $r_{ij}=|i-j|$ for $|i-j|\leqslant L/2$ and $r_{ij}=L-|i-j|$ for |ij|>L/2|i-j|>L/2, ${\cal N}=\sum_{x=1}^{L/2}\frac{1}{x^{\alpha}}$ is the normalization constant, and $\alpha$ is a parameter that controls the interaction range of the two-qubit gates.
+We simulate circuit dynamics using the tableau algorithm [Aaronson-PRA2004] (see also Supplemental Material [suppl]), for system sizes ranging from $L=16$ to $2048$, and evolution times up to $t_{\rm max}=4096$.
+The parameters $\alpha$ and $p$ are varied in the ranges $1\leqslant\alpha\leqslant 4$ and $0\leqslant p<1$, respectively.
 II.2 Entanglement entropy and quantum conditional mutual information
@@
 To compute the entanglement entropy in stabilizer circuits, we follow the method described in Ref.
-[24] (see also Supplemental Material [23]).
+[Nahum-PRX2017] (see also Supplemental Material [suppl]).
 Note that when $p=1$, the entanglement entropy vanishes for all subsystems since the state $|\phi_{\zeta}(t)\rangle$ becomes a product state immediately after every measurement.
@@
 II.3 Entanglement entropy at criticality
-Motivated by the emergent conformal criticality observed in monitored Clifford and non-Clifford circuits with short-range interactions or hoppings [4, 2, 25, 26, 12, 13], we assume that the steady-state entanglement entropy $S_{X}=\frac{1}{N_{t}}\sum_{k=0}^{N_{t}-1}\langle\langle S_{X,\zeta}(t_{\rm min}+k\Delta_{S_{X},t})\rangle\rangle$ at criticality behaves as [27]
+Motivated by the emergent conformal criticality observed in monitored Clifford and non-Clifford circuits with short-range interactions or hoppings [Skinner-PRX2019, Li-PRB2019a, Fuji-PRB2020, Alberton-PRL2021, Minato-PRL2022, Block-PRL2022], we assume that the steady-state entanglement entropy $S_{X}=\frac{1}{N_{t}}\sum_{k=0}^{N_{t}-1}\langle\langle S_{X,\zeta}(t_{\rm min}+k\Delta_{S_{X},t})\rangle\rangle$ at criticality behaves as [Calabrese-JPAMT2009]
 $S_{X}\sim\frac{c}{3}\log_{2}{l_{X}}+c^{\prime}\quad(p=p_{\rm c}),$ (5)
-where $l_{X}\equiv\frac{L}{\pi}\sin\frac{\pi|X|}{L}$ is the chord length of subsystem $X$ with $|X|$ denoting the number of qubits in $X$, $p_{\rm c}$ the critical value of the measurement rate, $c$ a universal coefficient of the logarithmic term (often referred to as an effective central charge) [28], and $c^{\prime}$ a nonuniversal constant.
+where $l_{X}\equiv\frac{L}{\pi}\sin\frac{\pi|X|}{L}$ is the chord length of subsystem $X$ with $|X|$ denoting the number of qubits in $X$, $p_{\rm c}$ the critical value of the measurement rate, $c$ a universal coefficient of the logarithmic term (often referred to as an effective central charge) [Jian-PRB2020], and $c^{\prime}$ a nonuniversal constant.
 The autocorrelation time $\Delta_{S_{X},t}$ for the steady-state entanglement entropy is estimated in the same manner as $\Delta_{I,t}$, and exhibits similar behavior.
 Notice that the coefficient $c$ here is not equivalent to the central charge in conformal field theory.
-For example, in the case of the percolation universality class, the central charge is known to be zero [28, 29].
+For example, in the case of the percolation universality class, the central charge is known to be zero [Jian-PRB2020, Oshima-PRB2023].
 Note also that in the deep volume-law phase ($p\ll p_{c}$), the entanglement entropy scales extensively as $S_{X}\sim|X|$, while in the area-law phase ($p\gg p_{\rm c}$), it scales with the boundary, $S_{X}\sim|\partial X|$, where $|\partial X|$ denotes the number of qubits on the boundary of subsystem $X$.
-Figure 2: Two distinct partitionings of the $L$-qubit system into subsystems $A$ (green), $B$ (white), $C$ (red), and $\overline{A\cup B\cup C}$ (gray).
+Figure 2: Two distinct partitions of the $L$-qubit system into subsystems AA (green), $B$ (white), $C$ (red), and $\overline{A\cup B\cup C}$ (gray).
 Each circle represents a qubit, and the number of qubits in each segment is indicated.
@@
 We note that in the area-law phase $p\gg p_{\rm c}$, the QCMI behaves as $I(A:C|B)\sim 0$, regardless of the partitioning.
-Moreover, the inequality $I(A:C|B)\geqslant 0$ always holds due to the strong subadditivity of von Neumann entanglement entropy [30].
+Moreover, the inequality $I(A:C|B)\geqslant 0$ always holds due to the strong subadditivity of von Neumann entanglement entropy [Lieb-PRL1973].
 Thus, as long as the partitioning satisfies conditions (i)-(iii), the QCMI $I(A:C|B)$ serves as a meaningful probe of the measurement-induced entanglement phase transition.
@@
 (5), implying that $I(A:C|B)$ can detect the entanglement phase transition even in the absence of emergent conformal invariance.
-Specifically, we consider two types of partitioning for the same system, in a spirit of Ref. [31]. The first partitioning sets $|A|=|B|=|C|=L/4$ [see Fig. 2(a)], for which we find
+Specifically, we consider two types of partitioning for the same system, in a spirit of Ref. [Furukawa-PRL2009]. The first partitioning sets $|A|=|B|=|C|=L/4$ [see Fig. 2(a)], for which we find
 $I_{a}(A:C|B)\sim\frac{c}{3}\log_{2}{\frac{\sin^{2}{\frac{\pi}{2}}}{\sin{\frac{\pi}{4}}\sin{\frac{3\pi}{4}}}}=\frac{c}{3}\quad\quad(p=p_{\rm c}).$ (7)
@@
 $I_{b}(A:C|B)\sim\frac{c}{3}\log_{2}{\frac{\sin^{2}{\frac{9\pi}{16}}}{\sin{\frac{\pi}{4}}\sin{\frac{\pi}{8}}}}\approx 1.83\times\frac{c}{3}\quad\quad(p=p_{\rm c}).$ (8)
-Here, the subscripts $a$ and $b$ in $I_{a}$ and $I_{b}$ correspond to the partitionings illustrated in Figs.
+Here, the subscripts $a$ and $b$ in $I_{a}$ and $I_{b}$ correspond to the partitions illustrated in Figs.
 2(a) and 2(b), respectively.
@@
 In what follows, we analyze the critical properties of the system along this phase boundary.
-Additionally, we confirm that no area-law phase exists for $\alpha\lesssim 2$, which is consistent with the previous analytical arguments in the high-measurement-rate regime [13, 15].
+Additionally, we confirm that no area-law phase exists for $\alpha\lesssim 2$, which is consistent with the previous analytical arguments in the high-measurement-rate regime [Block-PRL2022, Sahu-PRB2022].
 Figure 3: Phase diagram in the thermodynamic limit.
@@
 At the critical point $p=p_{\mathrm{c}}$, the QCMI becomes independent of the system size [see Eqs.
-(7) and (8)], which suggests that the crossing-point analysis [32] can be effectively used to estimate both $p_{\mathrm{c}}$ and the correlation-length critical exponent $\nu$.
+(7) and (8)], which suggests that the crossing-point analysis [Luck-PRB1985] can be effectively used to estimate both $p_{\mathrm{c}}$ and the correlation-length critical exponent $\nu$.
 Indeed, as shown in Fig.
 4, the curves of $I_{\gamma}(p,L)$ for different $L$ appear to cross at a single point for sufficiently large system sizes.
-The estimated critical points $p_{\mathrm{c}}$ are nearly identical for both partitionings $\gamma=a$ and $b$, for each value of $\alpha$.
+The estimated critical points $p_{\mathrm{c}}$ are nearly identical for both partitions $\gamma=a$ and $b$, for each value of $\alpha$.
 This already demonstrates, at least in part, that the QCMI serves as a reliable probe to detect the critical point.
@@
 4(b) and 4(d)].
-Nevertheless, well-established methods for handling corrections to scaling have been developed and documented in the literature [33].
+Nevertheless, well-established methods for handling corrections to scaling have been developed and documented in the literature [Shao-S2016].
 Interestingly, we find that corrections to scaling becomes more subtle for larger $\alpha$.
@@
 8(a) and 9(a)] emerges in the vicinity of the critical point, suggesting that multiple correction terms may need to be considered.
-A similar issue was only recently resolved in the study of the Néel-paramagnetic quantum phase transition in dimerized Heisenberg models, where the known value of the correlation-length exponent $\nu$ from the expected O(3) universality class was crucial for the analysis [34].
+A similar issue was only recently resolved in the study of the Néel-paramagnetic quantum phase transition in dimerized Heisenberg models, where the known value of the correlation-length exponent $\nu$ from the expected O(3) universality class was crucial for the analysis [Ma-PRL2018].
 In our case, however, neither the critical exponent $\nu$ nor the universality class is known a priori, necessitating the use of an alternative approach, as described in the following.
-Figure 4: QCMI $I_{\gamma}(p,L)$ for different partitionings and interaction ranges.
+Figure 4: QCMI $I_{\gamma}(p,L)$ for different partitions and interaction ranges.
 Panels (a) and (b) show results for $\gamma=a$, while (c) and (d) correspond to $\gamma=b$.
@@
 Here, we adopt an empirical approach to estimate $p_{\mathrm{c}}$ and $\nu$ in the thermodynamic limit based on a data-collapse analysis.
-We employ the Bayesian scaling analysis method [35] to collapse the data of $I_{\gamma}(p,L)$, assuming the following scaling ansatz:
+We employ the Bayesian scaling analysis method [Harada-PRE2011] to collapse the data of $I_{\gamma}(p,L)$, assuming the following scaling ansatz:
 $I_{\gamma}(p,L)=f_{\gamma}(\Delta pL^{1/\nu}),$ (9)
@@
 Instead, we perform data-collapse fits over the system size range $L\in[L_{\mathrm{min}},L_{\mathrm{max}}]$ with $L_{\mathrm{max}}=2048$ fixed and $L_{\mathrm{min}}$ systematically varied.
-We then examine the asymptotic behavior of the estimated values of $p_{\mathrm{c}}$ and $\nu$ as $L_{\mathrm{min}}$ increases [36, 37].
+We then examine the asymptotic behavior of the estimated values of $p_{\mathrm{c}}$ and $\nu$ as $L_{\mathrm{min}}$ increases [Otsuka-PRX2016, Otsuka-PRB2020].
 An important aspect of this analysis is the estimation of uncertainties in the fitted parameters.
@@
 6 and 7.
-We extrapolate the results to the thermodynamic limit, i.e., $01/L_{\mathrm{min}}=$ 0, using a linear fit based on the four largest values of $L_{\mathrm{min}}$.
+We extrapolate the results to the thermodynamic limit, i.e., $1/L_{\mathrm{min}}=0$, using a linear fit based on the four largest values of $L_{\mathrm{min}}$.
 Although the use of a linear fit and the specific number of data points are empirical choices, we find that the estimated values of $p_{\mathrm{c}}$ and $\nu$ in the thermodynamic limit are consistent between $\gamma=a$ and $\gamma=b$ within two standard deviations.
 As summarized in Table 1, the results obtained for other values of $\alpha$ are overall consistent between the two different partitioning schemes, agreeing to at least four (three) significant digits for $p_{\mathrm{c}}$ ($\nu$) in most cases of $\alpha$.
-Additional analyses are provided in Supplemental Material [23].
+Additional analyses are provided in Supplemental Material [suppl].
 Figure 5: Data-collapse fits of the QCMI $I_{\gamma}(p,L)$.
@@
 Solid and dashed lines represent linear fits to the data for $\gamma=a$ and $\gamma=b$, respectively.
-The extrapolated values in the thermodynamic limit ($1/L_{\mathrm{min}}\to 0$) are indicated by solid symbols at $01/L_{\mathrm{min}}=$ 0.
+The extrapolated values in the thermodynamic limit ($1/L_{\mathrm{min}}\to 0$) are indicated by solid symbols at $1/L_{\mathrm{min}}=0$.
 Figure 7: Same as Fig. 6, but for $\alpha=2.5$.
@@
 $p_{\mathrm{c}}$ $\nu$
-$\alpha$ $I=I_{a}$ $I=I_{b}$ $I=I_{a}$ $I=I_{b}$
+$\alpha$ $I_{a} I=I$$_{b} I=I$$_{a} I=I$$I_{b}$
 4.0 0.20363(1) 0.203617(9) 1.255(4) 1.248(4)
@@
 These points are denoted as $p_{\gamma}^{\times}(L,2L)$ and $I_{\gamma}^{\times}(L,2L)$, respectively.
-A key difficult in this analysis is the nonmonotonic finite-size behavior of the crossing points, as clearly illustrated in Figs.
+A key difficulty in this analysis is the nonmonotonic finite-size behavior of the crossing points, as clearly illustrated in Figs.
 8(a) and 9(a).
@@
 $I_{\gamma}^{\times}(L,2L)=\frac{c}{3}+b_{1}L^{-\omega_{1}}+b_{2}L^{-\omega_{2}},$ (11)
-where $a_{i}$ and $b_{i}$ ($i$=1,2) are constants, and the correction exponents are ordered such that $\omega_{2}>\omega_{1}>0$.
+where $a_{i}$ and $b_{i}$ (ii=1,2) are constants, and the correction exponents are ordered such that $\omega_{2}>\omega_{1}>0$.
 Note that $I_{b}^{\times}(L,2L)$ is normalized by a constant factor ($\simeq 1.83$) so that its thermodynamic limit matches that of $I_{a}^{\times}(L,2L)$, i.e., $c/3$, [see Eqs.
 (7) and (8)].
-There are two possible scenarios for the origin of the second correction term [34].
+There are two possible scenarios for the origin of the second correction term [Ma-PRL2018].
 One possibility is that it arises from a second irrelevant scaling field, independent of the first.
@@
 Although the scaling forms in Eqs.
-(10) and (11) appear relatively simple, we find that fitting both correction terms with sufficient precision is quite challenging [34, 33].
+(10) and (11) appear relatively simple, we find that fitting both correction terms with sufficient precision is quite challenging [Ma-PRL2018, Shao-S2016].
 This difficulty persists even when we fix $p_{\mathrm{c}}$ and $\nu$ in Eq.
@@
 Figure 8: System-size dependence of the crossing points $p_{\gamma}^{\times}(L,2L)$ for (a) $\alpha=3.5$ and (b) $\alpha=2.5$.
-Open circles (triangles) represent the results for $\gamma=a$ ($b$), while solid ones at $01/L=$ 0 indicate the values of $p_{\mathrm{c}}$ obtained by the data-collapse analysis shown in Figs.
+Open circles (triangles) represent the results for $\gamma=a$ ($b$), while solid ones at $1/L=0$ indicate the values of $p_{\mathrm{c}}$ obtained by the data-collapse analysis shown in Figs.
 6 and 7, as well as in Table 1.
@@
 $\omega_{1}$ $c/3$
-$\alpha$ $I=I_{a}$ $I=I_{b}$ $I=I_{a}$ $I=I_{b}$
+$\alpha$ $I_{a} I=I$$_{b} I=I$$_{a} I=I$$I_{b}$
 4.0 0.94(5) 0.92(7) 1.053(2) 1.052(2)
@@
 10, we plot the correlation-length exponent $\nu$ estimated from the data-collapse analysis (see Table 1) as a function of $\alpha$.
-We confirm that for $\alpha\gtrsim 3$, our model belongs to the same universality class as that of the short-range interacting models [13, 15].
-In particular, our numerical estimates are in full agreement with recent results for the model with nearest-neighbor Clifford gates [19], and differ significantly from the value predicted by classical two-dimensional percolation theory, $\nu=4/3$.
+We confirm that for $\alpha\gtrsim 3$, our model belongs to the same universality class as that of the short-range interacting models [Block-PRL2022, Sahu-PRB2022].
+In particular, our numerical estimates are in full agreement with recent results for the model with nearest-neighbor Clifford gates [Sierant-PRB2022], and differ significantly from the value predicted by classical two-dimensional percolation theory, $\nu=4/3$.
 The existence of a universality class behavior for $\alpha\gtrsim 3$ is further supported by the coefficient $c$ of the logarithmic term in the entanglement entropy, as shown in Fig.
@@
 This deviation from emergent conformal criticality corroborates previous findings of non-conformal-field-theory criticality, as evidenced by the deviation of the dynamical critical exponent $z$ from unity in Ref.
-[13].
+[Block-PRL2022].
 It is worth noting, however, that the critical point $p_{c}$ and the exponent $\nu$ remain consistent between the two partitioning schemes for all values of $\alpha$ (see Table 1).
@@
 Figure 10: $\alpha$-dependence of the correlation-length critical exponent $\nu$.
-The dotted line with the gray shaded area represents the result of the nearest-neighbor model, $\nu=1.265(15)$ [19].
+The dotted line with the gray shaded area represents the result of the nearest-neighbor model, $\nu=1.265(15)$ [Sierant-PRB2022].
 The inset shows a magnified view near $\alpha=3.0$.
@@
 Now we compare the coefficient $c/3$ of the logarithmic term in the entanglement entropy with those obtained for monitored nearest-neighbor random Clifford circuits in Refs.
-[2, 11, 19].
+[Li-PRB2019a, Zabalo-PRB2020, Sierant-PRB2022].
 While our analysis uses the base-2 logarithm, previous studies employed the natural logarithm in defining this coefficient.
@@
 From our estimate of $c/3=1.053(2)$ for the shortest-range case $\alpha=4$ using the partitioning scheme $\gamma=a$ (see Table 2), we obtain $\tilde{c}=1.519(3)$.
-This value is noticeably smaller than those in the previous studies: $\tilde{c}=1.6$ with up to 512 qubits [2], $\tilde{c}=1.61(3)$ with up to 512 qubits [11], and $\tilde{c}=1.57(1)$ with up to 10240 qubits [19] for monitored nearest-neighbor random Clifford circuits.
+This value is noticeably smaller than those in the previous studies: $\tilde{c}=1.6$ with up to 512 qubits [Li-PRB2019a], $\tilde{c}=1.61(3)$ with up to 512 qubits [Zabalo-PRB2020], and $\tilde{c}=1.57(1)$ with up to 10240 qubits [Sierant-PRB2022] for monitored nearest-neighbor random Clifford circuits.
 Furthermore, we find that $c/3$ exhibits only minor variation across different partitioning schemes ($\gamma=a,b$) and for a range of $\alpha\gtrsim 3$ [see the inset of Fig.
@@
 We obtain $\tilde{c}=1.6126(7)$, $1.5484(9)$, and $1.5263(7)$ [$c/3=1.1175(5)$, $1.0730(6)$, and $1.0577(5)$] for $\alpha=3.0$, $3.5$, and $4.0$, respectively.
-These results indicate that $c/3$ obtained from the entanglement entropy approaches the value estimated from the QCMI from above as $\alpha$ increases (see Supplemental Material [23] for results at other values of $\alpha$).
+These results indicate that $c/3$ obtained from the entanglement entropy approaches the value estimated from the QCMI from above as $\alpha$ increases (see Supplemental Material [suppl] for results at other values of $\alpha$).
 Since finite-size effects are expected to be smaller at larger $\alpha$, the estimates of $\tilde{c}$ from the entanglement entropy at criticality are more reliable for larger $\alpha$ than for the smaller ones.
@@
 Indeed, the values of $\tilde{c}$ estimated via the QCMI exhibit much weaker dependence on $\alpha$ [see the inset of Fig.
-11 and Table 2] compared to those obtained from the entanglement entropy [23].
+11 and Table 2] compared to those obtained from the entanglement entropy [suppl].
 We therefore adopt $\tilde{c}=1.519(3)$, obtained from the QCMI for $\alpha=4.0$, as our final estimate for the shortest-range interacting case.